OpenStudy (frostbite):

Anyone sitting on the original article by A. Fersht and the protein scientific conventions regarding protein stability?

OpenStudy (anonymous):

does it have a pubmed ID? if so, can you post it? I will see whether my institute has access.

OpenStudy (frostbite):

I wish I did, because in that case I could find it my self. The interest for this is due to a text I've been reading that talks about following the Fersht conventions. After reading the text I got to the conclusion that the text below could perhaps be the Fersht conventions, but I need the original article to see if this is actually true. Unfolding: \[\Large \sf N \leftrightarrow U\]\[\Large K_{eq}=\frac{ \left[ U \right] }{ \left[ N \right] }\]\[\Large \Delta G_{U-N}=-RT \ln(K_{eq})\]\[\Large \Delta G_{U-N}>0\] Folding: \[\Large \sf U \leftrightarrow N\]\[\Large K _{eq} = \frac{ \left[ N \right] }{ \left[ U \right] }\]\[\Large \Delta G _{N-U}=-RT \ln(K _{eq})\]\[\Large \Delta G_{N-U}<0\] Stability by mutagenesis: \[\large \Delta \Delta G_{U-N,MUT-WT}=\Delta G_{U-N,WT}-\Delta G_{U-N,MUT}\] \[\Large \Delta \Delta G_{U-N, MUT-WT}>0 ~ ~ \textrm{Destabilizing}\] \[\Large \Delta \Delta G_{U-N, MUT-WT}<0 ~ ~ \textrm{Stabilizing}\]

OpenStudy (frostbite):

This is indeed the way we calculate protein stability, and could require a convention, as there are no natural way of deriving them for a system without making the assumptions first as done above.

OpenStudy (frostbite):

found by part. The more detailed definitions are found in the references to the article linked in the question.

OpenStudy (anonymous):

wow. 2004!? I would've expected this work to be far older, actually. it's pretty comforting to know that ground work on physical chemistry is still a rewarding field to venture in =)

OpenStudy (frostbite):

I do think it is much older. Fersht's work goes back to 1989 which might be where he developed these definitions, but especially the mutagenesis stability was important as it is also included into \(\phi\)-analysis.

OpenStudy (frostbite):

I dare to say we still got the whole challenge with intrinsically disordered proteins. It most be governed by the laws of physical chemistry, but the relation may just not have been found yet,

OpenStudy (frostbite):

Even more I dare to say Fersht is THE pioneer in protein folding.